# A Study on Chinese Social Perspective regarding ChatGPT for Education and Beyond

### v3

@@ IV-B AI and Human
Negative |
As an AI chatbot, ChatGPT has almost become synonymous with “knowing everything”, but it can also spout nonsense. In reality, it simply re-edits known information, and its sources are unreliable. #Will ChatGPT replace humans?# |
Neutral |
#ChatGPT# AI is very intelligent, but AI is still AI. |
Positive |
Working with ChatGPT is convenient. For tasks like writing procedural emails, it makes sense to have AI do it! |
After the release of GPT-4, the sentiment analysis showed a significant decrease in the percentage of neutral sentiment, which is also supported by the ANOVA analysis showed in TableIV with an F-statistic of 106.6 and a p-value of 8.8e-37. Apparently, the launch of GPT-4, with its multi-modal capabilities enhancing applications in fields like image analysis and web design, has heightened people’s comprehension and knowledge of its abilities, inciting them to express positive or negative views on social platforms like Weibo. This suggests that as ChatGPT evolves and integrates more into daily life, awareness of its uses, risks, and dangers will grow, leading to a broader spectrum of attitudes towards AI technology.
@@
The digital age has heightened the importance of human-computer collaboration, especially with AI advancements. AI technologies like ChatGPT now understand and process human language, enhancing insights and assistance. This synergy enables novel approaches to complex tasks, enhancing efficiency, productivity, and accuracy in sectors like healthcare, finance, and education. Thus, we searched for relevant content on Weibo using keywords “human-computer collaboration” and “human-computer interaction” to gather insights and discussions relevant to these advancements.
As shown in Figure 7b and Table IV, there is no significant difference in the percentages after GPT-4’s launch, which is quite distinct from other scenarios. This may imply that most individuals do believe that ChatGPT can, to some degree, enhance human-computer collaboration and interaction, rather than being as pessimistic as its application in the field of education, where originality is of utmost importance. In other words, people are more positive of adopting GPTs in the areas in which originality is less important.
@@
Negative |
Fortunately, at present, AI is unable to write code; otherwise, I would be concerned about employment security. |
Neutral |
In the heated discussions around human-computer interaction, the focus of ChatGPT is not actually on the anxiety of whether it will replace humans or not, but on how humans can update their concepts and find the path to the future. |
Positive |
I quickly recognized the positive changes and enhancements that these new AI products have brought to my work. They have saved me substantial time in creating documents, Copilot may even automate the generation of my PPTs in the future. |
@@
People have consistently expressed concerns about the ethical and legal impact of AI. As shown in Figure 7d and Table IV, attitudes towards such AI tools about ethics and law remain remarkably consistent both before and after the launch of GPT-4, with a mere 1.3 and 1.2 having a comprehensive positive view, which is the lowest among all the scenarios. Concerns exist about ChatGPT’s ability to breach privacy, produce harmful responses, encourage unethical content, and potentially incite criminal behavior, thus undermining laws and moral values. Nonetheless, some argue that what humans should fear is becoming the Satan of the earth, not the world under AI. From their perspective, ChatGPT could serve as a valuable tool, driving positive societal developments.
@@
Negative |
I have cracked ChatGPT’s restrictions on ethical answers, and I will not disclose how to crack it because if it is used by bad people, the world will be in danger. |
Neutral |
It is the general trend for AI to replace technical jobs, but in fields that challenge the basic survival of human beings, it is impossible to completely hand over to AI. |
Positive |
#ChatGPT#It is crucial to foster a gradual understanding and active participation in this transformation while conscientiously considering the ethical, legal, and philosophical implications that emerge from the convergence of these new technologies. |
@@
Table X shows some people have cracked the restrictions of ChatGPT on generating ethical answers, and worry that this method may be abused and lead to harm. In fact, OpenAI has been fine-tuning the model based on human feedback [35], and the probability of InstructGPT generating toxic content is considerably lower than that of GPT-3. It is worthwhile to further fine-tune ChatGPT to better conform to ethics and morality so that it can become a tool to benefit humanity.
@@ V Discussion
Many new technologies, including ChatGPT, often lead to widespread discussions and emotional reactions on social media. Qaiser et al. [37] showed that people generally have negative feelings about the impact of technology on employment and advancements in AI. In educational settings, Clarizia et al. [38] found that students initially felt disoriented when introduced to new technologies, but their sentiments turned positive after teaching styles were adjusted to their emotional needs and more content was introduced. When introducing innovative products like Google Glass, people’s initial reaction was often surprise, a typically neutral emotion that can potentially shift toward negative feelings [39]. This negativity bias, where negative reactions outweigh positive ones, could explain why some technological products fail. Thus, guiding public emotion positively is vital during the launch and use of new technologies. Based on our research results, we should adopt certain strategies like leveraging social network to promote positive user feedback to shift public sentiment about ChatGPT from negative to positive, ensuring its reasonable application in education and other fields.
@@
The comparison between our study on Weibo and a related study by Tlili et al. [15] on X reveals some similarities and differences in the trends observed. One of the similarities is that neutral or unclassified sentiment dominates the majority of the comments related to ChatGPT’s application in education, reflecting the complex and nuanced nature of people’s attitudes towards this technology. Another similarity is that both studies highlight some common advantages and concerns regarding ChatGPT’s use in education, indicating consistency in people’s perceptions across social media platforms.
@@
However, there are also some notable differences between the two studies. Our study shows that negative sentiment represents a significantly higher proportion than positive sentiment towards ChatGPT’s application in education, in contrast to the related study that reports a higher proportion of positive sentiment. We analyzed the posts and identified several key factors contributing to it. Firstly, unlike in the West, users often require a VPN to access ChatGPT in China, and some service providers even charge for ChatGPT-3.5 that is typically free., which significantly diminishes the user experience. Secondly, concerns regarding ChatGPT’s security have been raised by various organizations and individuals in China, resulting in a preference for domestic LLMs over international ones. These factors collectively play a significant role in shaping the overall sentiment towards ChatGPT among Chinese users.
@@
In contrast to its educational applications, ChatGPT’s use in sectors like industry, involving human-computer interaction, yields similar insights from X [40] and Weibo data. As highlighted by Aram Bahrini et al. [18], while ChatGPT boosts efficiency across various scenarios, it’s vital to consider ethical, privacy, and security concerns, along with the risk of errors and the need for inclusiveness. This echoes our analysis’s second part, acknowledging ChatGPT’s benefits, like enhanced productivity and cost-saving potential, alongside the importance of being aware of associated risks.
@@
When it comes to biases, there is a concept that is gaining increasing attention: Inclusive AI [23, 24, 25], , which means AI should be able to showcase the world’s diverse cultures and backgrounds, yet human-created AI often fails to accurately reflect this richness due to various limitations. Therefore, exploring strategies to enhance fairness, inclusivity, and user-friendliness in human-created AI is essential. For instance, Druga et al.’s study [41] revealed that factors like nationality, age, and socio-economic status influence children’s engagement with AI, highlighting the need for expert input. Ovalle et al. [42] found that transgender and non-binary individuals face heightened discrimination and exclusion. Their research focused on evaluating the impact of misgendering and negative responses to gender disclosure in open language generation, exposing the dominance of binary gender norms and underscoring the urgent need for more inclusive AI approaches, including in LLMs and other models.
@@ VI Conclusion
This study undertakes a comprehensive social survey focusing specifically on China’s landscape to delve into public opinions regarding the application of ChatGPT in education and other fields such as human-computer collaboration. The research reveals that most people maintain a neutral or negative attitude towards those applications due to concerns such as personal privacy infringement, content bias, and the ruin of originality, particularly in education. However, with the release of GPT-4, more people are recognizing its potential, resulting in a higher proportion of positive attitudes. Nevertheless, the majority still maintains a negative perception of AI, with concerns raised regarding its potential problems. In light of these findings, this study emphasizes the importance of considering public perceptions and concerns when developing and implementing AI technologies like ChatGPT in various fields. It is crucial to alleviate worries to ensure the successful adoption of ChatGPT while minimizing potential risks.

### v4

@@ IV-B AI and Human
| Sentiment | Artificial Intelligence | Human-Computer Collaboration and Interaction | Law, Morals and Ethics | |||
| p | p | p | ||||
| Positive | ||||||
| Neutral | ||||||
| Negative | ||||||
@@
| Negative | As an AI chatbot, ChatGPT has almost become synonymous with “knowing everything”, but it can also spout nonsense. In reality, it simply re-edits known information, and its sources are unreliable. #Will ChatGPT replace humans?# |
| Neutral | #ChatGPT# AI is very intelligent, but AI is still AI. |
| Positive | Working with ChatGPT is convenient. For tasks like writing procedural emails, it makes sense to have AI do it! |
After the release of GPT-4, the sentiment analysis showed a significant decrease in the percentage of neutral sentiment, which is also supported by the chi-square test results showed in TableVIII with and . Apparently, the launch of GPT-4, with its multi-modal capabilities enhancing applications in fields like image analysis and web design, has heightened people’s comprehension and knowledge of its abilities, inciting them to express positive or negative views on social platforms like Weibo. This suggests that as ChatGPT evolves and integrates more into daily life, awareness of its uses, risks, and dangers will grow, leading to a broader spectrum of attitudes towards AI technology.
@@
The digital age has heightened the importance of human-computer collaboration, especially with AI advancements. AI technologies like ChatGPT now understand and process human language, enhancing insights and assistance. This synergy enables novel approaches to complex tasks, enhancing efficiency, productivity, and accuracy in sectors like healthcare, finance, and education. Thus, we searched for relevant content on Weibo using the keywords “human-computer collaboration” and “human-computer interaction” to gather insights and discussions relevant to these advancements.
As shown in Figure 7b and Table VIII, there is no significant difference in the percentages after GPT-4’s launch, which is quite distinct from other scenarios. This may imply that most individuals do believe that ChatGPT can, to some degree, enhance human-computer collaboration and interaction, rather than being as pessimistic as its application in the field of education, where originality is of utmost importance. In other words, people are more positive of adopting GPTs in the areas in which originality is less important.
@@
| Negative | Fortunately, at present, AI is unable to write code; otherwise, I would be concerned about employment security. |
| Neutral | In the heated discussions around human-computer interaction, the focus of ChatGPT is not actually on the anxiety of whether it will replace humans or not, but on how humans can update their concepts and find the path to the future. |
| Positive | I quickly recognized the positive changes and enhancements that these new AI products have brought to my work. They have saved me substantial time in creating documents, Copilot may even automate the generation of my PPTs in the future. |
@@
People have consistently expressed concerns about the ethical and legal impact of AI. As shown in Figure 7d and Table VIII, attitudes towards such AI tools about ethics and law remain remarkably consistent both before and after the launch of GPT-4, with a mere 1.3 and 1.2 having a comprehensive positive view, which is the lowest among all the scenarios. Concerns exist about ChatGPT’s ability to breach privacy, produce harmful responses, encourage unethical content, and potentially incite criminal behaviour, thus undermining laws and moral values. Nonetheless, some argue that what humans should fear is becoming the Satan of the earth, not the world under AI. From their perspective, ChatGPT could serve as a valuable tool, driving positive societal developments.
@@
| Negative | I have cracked ChatGPT’s restrictions on ethical answers, and I will not disclose how to crack it because if it is used by bad people, the world will be in danger. |
| Neutral | It is the general trend for AI to replace technical jobs, but in fields that challenge the basic survival of human beings, it is impossible to completely hand over to AI. |
| Positive | #ChatGPT#It is crucial to foster a gradual understanding and active participation in this transformation while conscientiously considering the ethical, legal, and philosophical implications that emerge from the convergence of these new technologies. |
@@
Table XI shows some people have cracked the restrictions of ChatGPT on generating ethical answers, and worry that this method may be abused and lead to harm. In fact, OpenAI has been fine-tuning the model based on human feedback [35], and the probability of InstructGPT generating toxic content is considerably lower than that of GPT-3. It is worthwhile to further fine-tune ChatGPT to better conform to ethics and morality so that it can become a tool to benefit humanity.
@@ V Discussion
Many new technologies, including ChatGPT, often lead to widespread discussions and emotional reactions on social media. Qaiser et al. [37] showed that people generally have negative feelings about the impact of technology on employment and advancements in AI. In educational settings, Clarizia et al. [38] found that students initially felt disoriented when introduced to new technologies, but their sentiments turned positive after teaching styles were adjusted to their emotional needs and more content was introduced. When introducing innovative products like Google Glass, people’s initial reaction was often surprised, a typically neutral emotion that can potentially shift toward negative feelings [39]. This negativity bias, where negative reactions outweigh positive ones, could explain why some technological products fail. Thus, guiding public emotion positively is vital during the launch and use of new technologies.
Based on our research results, we should adopt certain strategies like leveraging social networks to promote positive user feedback to shift public sentiment about ChatGPT from negative to positive, ensuring its reasonable application in education and other fields. Meanwhile, it is crucial to train teaching staff to improve their digital literacy, ensuring they are aware of the limitations of AI tools, use them responsibly, and guide students in critical thinking, which aligns with the views of other researchers [40, 41, 42, 43, 44].
@@
The comparison between our study on Weibo and a related study by Tlili et al. [15] on reveals some similarities and differences in the trends observed. One of the similarities is that neutral or unclassified sentiment dominates the majority of the comments related to ChatGPT’s application in education, reflecting the complex and nuanced nature of people’s attitudes towards this technology. Another similarity is that both studies highlight some common advantages and concerns regarding ChatGPT’s use in education, indicating consistency in people’s perceptions across social media platforms.
@@
However, there are also some notable differences between the two studies. Our study shows that negative sentiment represents a significantly higher proportion than positive sentiment towards ChatGPT’s application in education, in contrast to the related study that reports a higher proportion of positive sentiment. We analyzed the posts and identified several key factors contributing to it. Firstly, compared to other LLMs like Ernie Bot, although ChatGPT has received positive feedback and enthusiasm from Chinese users for its superior performance, there is currently no straightforward method to use ChatGPT within the mainland, which impacts the user experience. Additionally, ChatGPT’s responses sometimes display misunderstandings or biases about Chinese culture and society. Users sometimes observe ChatGPT’s understanding and generation of Chinese are less natural and accurate. These factors collectively play a significant role in shaping the overall sentiment towards ChatGPT among Chinese users.
@@
In contrast to its educational applications, ChatGPT’s use in sectors like industry, involving human-computer interaction, yields similar insights from [45] and Weibo data. As highlighted by Aram Bahrini et al. [18], while ChatGPT boosts efficiency across various scenarios, it’s vital to consider ethical, privacy, and security concerns, along with the risk of errors and the need for inclusiveness. This echoes our analysis’s second part, acknowledging ChatGPT’s benefits, like enhanced productivity and cost-saving potential, alongside the importance of being aware of associated risks.
@@
When it comes to biases, there is a concept that is gaining increasing attention: Inclusive AI [23, 24, 25, 46], which means AI should be able to showcase the world’s diverse cultures and backgrounds, yet human-created AI often fails to accurately reflect this richness due to various limitations. Therefore, exploring strategies to enhance fairness, inclusivity, and user-friendliness in human-created AI is essential. For instance, Druga et al.’s study [47] revealed that factors like nationality, age, and socio-economic status influence children’s engagement with AI, highlighting the need for expert input. Ovalle et al. [48] found that transgender and non-binary individuals face heightened discrimination and exclusion. Their research focused on evaluating the impact of misgendering and negative responses to gender disclosure in open language generation, exposing the dominance of binary gender norms and underscoring the urgent need for more inclusive AI approaches, including in LLMs and other models.
@@ VI Conclusion
This study undertakes a comprehensive social survey focusing specifically on China’s landscape to delve into public opinions regarding AI applications such as ChatGPT in education and human-computer collaboration and interaction. The research reveals that most people maintain a neutral or negative attitude towards those applications due to concerns such as personal privacy infringement, content bias, and the ruin of originality, particularly in education. However, as AI tools evolve, more people are recognizing their potential, resulting in a higher proportion of positive attitudes. Nevertheless, the majority still maintains a negative perception of AI, with concerns raised regarding its potential problems. In light of these findings, this study emphasizes the importance of considering public perceptions and concerns when developing and implementing AI technologies like ChatGPT in various fields. It is crucial to alleviate worries to ensure the successful adoption of AGI while minimizing potential risks.